OpenStudy (<3pandas18):

A fox sees a piece of carrion being thrown from a hawk’s nest and rushes to snatch it. The nest is 14.0 m high, and the carrion is thrown with a horizontal velocity of 1.5 m/s. The fox is 7.0 m from the base of the tree. What is the magnitude of the fox’s average velocity if it grabs the carrion in its mouth just as it touches the ground? (Neglect air resistance; a = g = 9.81 m/s.)

563blackghost (563blackghost):

This is for the quiz in projectile motion right?

OpenStudy (joseph127):

Well this is a quit easy question if you have an image yo illustrate it

OpenStudy (joseph127):

Okay now lets solve the question H= height of tree g=g-force t=time S=distance u=Carrion's velocity V=fox velocity Okay now we first find how long it took the carrion to fall 1). H=1/2*g*t^2 14=0.5*9.8*t^2 t=1.7seconds Now that we know how long it took to fall we now solve for the position on the x-axis it fell on 2). t=s/u 1.7=s/1.5 s=2.55m Now that we know its position on the x-axis we now find how fast the fox has to move to catch the carrion in its mouth s=7-2.55. s=4.45m V=s/t V=4.45/1.7 V=2.62m/s So the fox has to move at a velocity of 2.62m/s to be able to catch the carrion as soon as it gets to ground level. Hope am right

563blackghost (563blackghost):

Agreed.